Solve the system by the elimination method.
-4x=25-3y
6x=5y-39

Answer by mananth(16946)

-4 x + 3 y = 25 .............1
Total value
6 x -5 y = -39 .............2
Eliminate y
multiply (1)by 5
Multiply (2) by 3
-20 x 15 y = 125
18 x -15 y = -117
Add the two equations
-2 x = 8
/ -2
x = -4
plug value of x in (1)
-4 x + 3 y = 25
16 + 3 y = 25
3 y = 25 -16
3 y = 9
y = 3
x= -4
y= 3
m.ananth@hotmail.ca